OpenStudy (anonymous):

An airplane flying at a speed of 310 mi/hr flies from a point A in the direction 132° for 30 minutes and then flies in the direction 222° for 45 minutes. Approximate, to the nearest mile, the distance from the airplane to A.

OpenStudy (anonymous):

 you can see this is a right angles triangle. we want d3 but first we need either d1 or d2. lets find d1. velocity=distance/time so v=d1/t1, d1=vt1, d1= 310*0.5=155 miles. Now use right angles trig to find d3. use costheta=d1/d3, ive placed most of the angles there which i worked out using simple trig, the angle you need work it out using the fact that the sum of angles inside a triangle= 180 degrees. :D